# Generative AI-enabled Wireless Communications for Robust Low-Altitude Economy Networking

### v1

@@ V-D Numerical Results
In the experiments, we consider a terrestrial BS with a $4\times 4$ UPA, a legitimate UAV, and an eavesdropper, each equipped with $N_{b}=N_{e}=6$ antennas. The beamforming design accounts for both position and channel uncertainties while adopting the same channel power gain settings as in [14], and $C_{eve}=3b/s/Hz$. In chance-constrained optimization, we set the constraint probability $P_{eve}=70\%$. We evaluate our method benchmarking with the SAC [15] and GDM [13].
The learning curve for robust optimization is presented in Fig. 4(a), where the red, green, and blue curves represent the MoE-Transformer, GDM, and SAC algorithms, respectively. Moreover, the dashed lines are the average ASR reward under robust optimization. Under the same number of training epochs, the convergence reward of our algorithm is the highest. For stochastic and chance-constrained optimization, our proposed algorithm achieves over 200% improvement compared to GDM. Compared to the SAC algorithm, our model exhibits an improvement of 27.93% and 19.37%, respectively. In the robust optimization, the GDM algorithm outperforms the SAC algorithm, indicating that the GenAI-enabled model enhances system robustness and improves its ability to handle uncertainties. Nonetheless, our proposed model still achieves the best performance, showing improvements of 44.33% and 58.70% over the two benchmark methods. These results further highlight how the MoE-transformer architecture significantly strengthens system robustness.
Finally, we compared the normalized inference reward of our proposed method. Across the three optimization paradigms—stochastic, chance-constrained, and robust optimization—each algorithm achieves the best performance or suboptimal under its own optimization target. However, robust optimization exhibits the weakest performance for the other two criteria, while the chance-constrained approach maintains a more balanced performance. This behavior can be attributed to the inherent fading characteristics of wireless channels, such as the Rayleigh fading model. The robustness optimization strategy is not effective during an outage, where we have to allocate infinite power for channel compensation. In contrast, chance-constrained optimization effectively mitigates extreme values in a probabilistic manner, resulting in better overall performance. Therefore, in practical scenarios, absolute robustness (i.e., optimizing for the worst-case in every situation) may not be the most effective approach. Instead, incorporating probabilistic error tolerance allows for the design of robust LAENets that better align with real-world requirements.
VI Conclusion
@@
In this paper, we have explored the application of GenAI techniques to enhance the robustness of LAENets. Initially, we conducted a systematic analysis of the robustness requirements for LAENets, followed by a comprehensive review of robust QoS metrics from the perspective of the wireless physical layer. Subsequently, we have performed an in-depth analysis of existing GenAI-enabled approaches for robustness enhancement and their potential applications. Building on this, we have proposed a diffusion-based reinforcement framework with an MoE-transformer actor network. In the robust beamforming case study, the proposed framework has effectively learned the beam pattern with various uncertainties, achieving an improvement of approximately 19% ASR in stochastic and chance-constrained optimization, and around 44% ASR in robust optimization. These results underscore the critical role of GenAI in enhancing the robustness of LAENets and highlight the need for further exploration of its applications.
Biographies
Changyuan Zhao is a PhD candidate at Nanyang Technological University, Singapore.
Jiacheng Wang is a research fellow at Nanyang Technological University, Singapore.
Ruichen Zhang is a research fellow at Nanyang Technological University, Singapore.
Dusit Niyato is a chair professor with Nanyang Technological University, Singapore.
Geng Sun is a professor with Jilin University, China.
Hongyang Du is an assistant professor with The University of Hong Kong, Hong Kong.
Dong In Kim is a professor with Sungkyunkwan University, South Korea.
Abbas Jamalipour is a professor with University of Sydney, Australia.

### v2

@@ V-D Numerical Results
In the experiments, we consider a terrestrial BS with a $\times 4$ 4 UPA, a legitimate UAV, and an eavesdropper, each equipped with $N_{b}=N_{e}=6$ antennas. The beamforming design accounts for both position and channel uncertainties while adopting the same channel power gain settings as in [13], and $C_{eve}=3bps/Hz$. In chance-constrained optimization, we set the constraint probability $P_{eve}=70\%$. We evaluate our method by benchmarking it against the non-GenAI method SAC [14], GNN [15], transformer [12], and the GenAI method GDM [4], which does not use an MoE-transformer Actor Network. All networks are trained for 2000 epochs with a learning rate of $1\times 10^{-4}$. Our MoE-transformer actor network [^4] For more implementation details, please refer to the code available at https://github.com/ChangyuanZhao/LAENet_robust_opt. employs 4 experts with top-2 selection, uses 4 attention heads, and adopts a hidden dimension of 256. In the GDM setting, we configure the diffusion process with 6 steps and adopt the same network architecture as in [4].
The learning curves for the three optimizations are presented in Figs. 4(a), 4(b), and 4(c), where the red, green, blue, purple, and cyan curves represent the MoE-Transformer, GDM, SAC, GNN, and transformer algorithms, respectively. Moreover, the dashed lines in Fig. 4(c) are the average ASR reward under robust optimization. The results show that our proposed algorithm consistently demonstrates the best learning capability across all optimization settings. In stochastic optimization, it moves beyond SAC and GNN by about 13%. Compared to GDM algorithm, whose learning curve remains noisy and slow to settle, our proposed algorithm achieves over 200% improvement. In chance-constrained optimization, most algorithms exhibit similar convergence behavior to our proposed method, which maintains a relatively fast convergence speed. In robust optimization, where all returns shift into negative territory, the MoE-transformer remains the least degraded. Its reward sits around 15.8% above transformer and about 44% higher than SAC. Collectively, these curves show that pairing a sparse mixture-of-experts backbone with diffusion-guided policy learning not only accelerates convergence but also provides consistently superior robustness across the full spectrum of optimization objectives. Although the MoE-Transformer activates 4 experts, its per-iteration wall-clock time remains 0.01906 s, versus 0.01813 s for the transformer, 0.01625 s for SAC and GDM, and 0.01875 s for GNN. This around 15% latency gap is far below the 100 ms control-loop budget typical for LAE edge nodes. Consequently, the proposed architecture satisfies real-time and energy-efficiency requirements while still delivering the robustness gains reported above.
Finally, we compared the inference reward of all methods in robust optimization as shown in Fig. 4(d). The proposed MoE-transformer method achieves the lowest negative minimum reward with the lowest variance of 16.99, indicating significantly steadier behavior. In contrast, GDM, SAC, GNN, and the transformer model all show larger variances, indicating that their worst-case returns fluctuate more significantly, and occasionally result in extremely low returns. The less variance seen in the MoE-transformer’s box plot results in a more dependable safety margin against adverse conditions, with more robust beamforming policies.
V-E Open Challenges
Despite the encouraging gains shown in Figs. 4(a)4(d), our study also uncovers several open challenges that must be tackled before GenAI-enabled robust beamforming can be translated from the lab to large-scale LAENets.
  • Data requirement: GenAI needs to learn the policy distributions. Greater uncertainties directly translate to a larger data requirement before the optimal LAENet strategy can be identified.
  • Hyper-parameter sensitivity: Performance depends on the noise schedule, denoising depth, and MoE expert count. Field deployments will demand continual re-tuning to balance latency, energy, and robustness.
  • Monte-Carlo simulation limitation: Current solutions justify reliability with empirical simulations only. Without formal worst-case guarantees, true end-to-end robustness remains unproven for safety-critical aerial links.
VI Future Directions
VI-A Adaptive GenAI for Dynamic Environments
LAENets, such as drone-based logistics and aerial monitoring systems, operate in highly dynamic and unpredictable environments. Customizing GenAI to adapt to real-time changes in network topology, environmental conditions, and mission requirements is essential. Techniques, including online learning, real-time model adaptation, and context-aware optimization can be explored to ensure that GenAI remains robust and effective in these ever-changing scenarios.
VI-B GenAI-enabled Anomaly Detection and Recovery
Ensuring reliability in LAENets requires robust mechanisms for detecting and recovering from anomalies such as signal interference, equipment malfunctions, and unexpected obstacles. GenAI can be utilized to develop advanced anomaly detection frameworks that integrate multimodal data, including sensor, visual, and communication data. Furthermore, designing GenAI models that suggest real-time recovery strategies, such as dynamic path adjustments or alternate resource allocations, can significantly enhance network resilience.
VI-C Energy-Aware GenAI Optimization
Energy efficiency is a critical constraint in LAENets, where nodes often rely on limited power sources such as drone batteries. Future research should focus on developing energy-aware GenAI models that balance computational demands with power constraints. Techniques such as lightweight model architectures, energy-efficient inference, and collaborative processing across network nodes can help extend operational lifespans while maintaining the robustness of the network.
VII Conclusion
@@
In this paper, we have explored the application of GenAI techniques to enhance the robustness of LAENets. Initially, we conducted a systematic analysis of the robustness requirements for LAENets, followed by a comprehensive review of robust QoS metrics from the perspective of the wireless physical layer. Subsequently, we have performed an in-depth analysis of existing GenAI-enabled approaches for robustness enhancement and their potential applications. Building on this, we have proposed a diffusion-based reinforcement framework with an MoE-transformer actor network. In the robust beamforming case study, the proposed framework has effectively learned the beam pattern with various uncertainties, achieving an improvement of approximately 13% ASR in stochastic optimization, and around 15% ASR in robust optimization than four learning baselines. These results underscore the critical role of GenAI in enhancing the robustness of LAENets and highlight the need for further exploration of its applications.